# Revisiting Non-Rotating Star Models: Classical Existence and Uniqueness Theory and Scaling Relations

### v1

@@ 2.2 Existence Results for Non-rotating Bodies
$=f(\pi)$
When $\pi$ is large enough we have $f(\pi)\leq-l$. Therefore, the claim holds true. Thus $\forall\pi>\pi_{0},\forall x$ with $r(x)>\pi$, (notice $r_{0}\in\left(1,\frac{1+\pi}{2}\right)$) we have $V_{\sigma_{R}}(x)\leq-l,\sigma_{R}(x)=0$.
Now we consider the z-direction and show $V_{\sigma_{R}}(x)\leq-l$ holds when $z(x)$ is large enough, which is relatively easier. We can first show $\exists\delta>0,r\geq 2$ and $\widetilde{R_{0}}>0$, such that $\forall R\geq\widetilde{R_{0}}$, if $\int_{|z(x)-d|<r}\sigma_{R}(x)d^{3}x<\delta$, then $\sigma_{R}(x)=0$ for $|z(x)-d|\leq 1$. This comes from that when $|z(x)-d|\leq 1$ we have the estimate:
$V_{\sigma_{R}}(x)=\int_{|z(y)-d|<r}\frac{\sigma_{R}(x)}{|y-x|}d^{3}y+\int_{|z(y)-d|>r}\frac{\sigma_{R}(x)}{|y-x|}d^{3}y\leq c\left(\delta^{a}+\delta^{b}\right)+\frac{m}{r-1}$
And then letting $\delta$ small enough and $r$ large enough can give $V_{\sigma_{R}}(x)\leq-l$, which means $\sigma_{R}(x)=0$. Then we set $Z_{n}=\{x||z(x)-2n\mid\leq 1\},n=0,\pm 1,\pm 2,\ldots$. and $Z_{n}^{\prime}=\{x||z(x)-2n\mid<r\}$, without loss of generality, we can take $r$ to be an integer. Let $N$ be the number of $Z_{n}^{\prime}$ such that $\int_{Z_{n}^{\prime}}\sigma_{R}\geq\delta$. Since the mass of $\sigma_{R}$ is $m<\infty$ and the $Z_{n}^{\prime}$, overlap finite times $\widetilde{r}$$\infty$, we have $N\delta\leq\widetilde{r}m$ or $N\leq\frac{\widetilde{r}m}{\delta}$. Therefore, $\sigma_{R}(x)=0$ for large $R$ except on a set of horizontal slabs $Z_{n}$ of total height at most $\frac{2\widetilde{r}m}{\delta}$.
If these $Z_{n}{s}$ were known to be contiguous, then we would have $\sigma_{R}(x)=0$ for $|z(x)|\geq\frac{2rm}{\delta}$. But if the $Z_{n}{s}$ were not contiguous, we could follow the idea of the strong rearrangement inequality in Lieb [14, Lemma 3]. We could rearrange the order of $Z_{n}$, slide the half-space above an empty $Z_{n}$ down one unit and obtain a new $\sigma_{R}^{\prime}\in W_{R}$ from $\sigma_{R}$. During the process we notice $U\left(\sigma_{R}^{\prime}\right)=\int A\left(\sigma_{R}^{\prime}\right)=\int A\left(\sigma_{R}\right)=U\left(\sigma_{R}\right),G\left(\sigma_{R}^{\prime},\sigma_{R}^{\prime}\right)=\int\sigma_{R}^{\prime}V_{\sigma_{R}^{\prime}}>\int\sigma_{R}V_{\sigma_{R}}=G\left(\sigma_{R},\sigma_{R}\right)$, then we obtain $E_{0}\left(\sigma_{R}^{\prime}\right)<E_{0}\left(\sigma_{R}\right)$, which leads to a contradiction.
@@
Thanks to Lemma 2.22 and Lemma 2.28, we know $\forall R\geq R^{\prime},\sigma_{R}\in W_{R^{\prime}}$, which implies $E_{0}\left(\sigma_{R}\right)=E_{0}\left(\sigma_{R^{\prime}}\right)$ since they should both be minimizers on $W_{R^{\prime}}$. Thus, we just need to show $\sigma_{R^{\prime}}$ is a minimizer on $W_{\infty}=m{R}\left(\mathbb{R}^{3}\right)$. To show that, given $\rho\in W_{\infty}$, we construct a $\rho_{R}\in W_{R}$ for large $R$, with $E_{0}\left(\rho_{R}\right)\leq E_{0}(\rho)+\epsilon$, where $\epsilon$ can be arbitrary small. (For the construction of $\rho_{R}$, although $\rho$ is not necessarily to be spherically symmetric, but one can apply a strong rearrangement inequality in Lieb [14, Lemma 3] and the generalization by Sobolev of the Hardy-Littlewood theorem on rearrangements of functions [21] to replace $\rho$ by another spherically symmetric and radially decreasing (after translation) $\rho$, see for example [1, Appendix], [16, Section 6]. Then one can refer to [1, Section 6] or [6, Theorem 3.1.7] to construct $\rho_{R}$.) Then we have $E_{0}\left(\sigma_{R^{\prime}}\right)=E_{0}\left(\sigma_{R}\right)\leq E_{0}\left(\rho_{R}\right)\leq E_{0}(\rho)+\epsilon$. Let $\epsilon\rightarrow 0$ we obtain $E_{0}\left(\sigma_{R^{\prime}}\right)\leq E_{0}(\rho)$, which shows the result. ∎
It is natural to conjecture if there are some relations between Lagrange-multipliers $\lambda_{m}$ and the derivatives of $e_{0}(m)$, since they can be obtained by somehow “differentiating” $E_{0}\left(\sigma_{m}\right)$. Thanks to Lieb and Yau’s paper [13, Theorem 3 and Theorem 4], we can know the left and right derivatives of $e_{0}(m)$ bound $\lambda:e_{0}{}^{\prime}\left(m^{+}\right)\leq\lambda\leq e_{0}{}^{\prime}\left(m^{-}\right)$, as cited in McCann’s paper [16, Theorem 3.5] and [7, Theorem 2.25 (viii)] above. Moreover, we find more precise quantitative property in the case that complies with polytropic equations of state in the following.
@@
By scaling relations in Section 3 (Theorem 3.2) we can obtain $e_{0}(m)=m^{\frac{5\gamma-6}{3\gamma-4}}e_{0}$, which implies $e_{0}^{\prime}$ exists. Then $e_{0}^{\prime}(m)=\lambda_{m}$ can come from Theorem 2.6 (viii). As in the proof of Theorem 2.6 (ii), we use the scaling relations mentioned in Section 3 (Proposition 3.4). We obtain $E_{0}^{\prime}\left(\sigma_{m}\right)(x)=\frac{1}{A^{\gamma-1}}E_{0}^{\prime}(\sigma)\left(\frac{1}{B}x\right)$, thus $\lambda_{m}=\frac{1}{A^{\gamma-1}}\lambda_{1}$. Since $e_{0}(m)=m^{\frac{5\gamma-6}{3\gamma-4}}e_{0}$, then $e_{0}^{\prime}(m)=\frac{5\gamma-6}{3\gamma-4}m^{\frac{2\gamma-2}{3\gamma-4}}e_{0}$, and $e_{0}^{\prime}(m)=\frac{1}{A^{\gamma-1}}e_{0}^{\prime}(1)$. Therefore, we just need to consider the case $m=1$. Notice we have Euler-Lagrange equation $E_{0}^{\prime}(\sigma)(x)=\lambda_{1}$ a.e. where $\sigma(x)>0$. Rather than substituting specific $x$ to calculate $E_{0}^{\prime}(\sigma)$, we view $E_{0}^{\prime}(\sigma)$ as a functional in the sense that $E_{0}^{\prime}(\sigma)(\rho)=\int E_{0}^{\prime}(\sigma)\rho$ as in Lemma 2.8 or Remark 2.9. We take $\rho=\sigma$ as a test function, notice the mass of $\sigma$ is 1. Then we have
$\int E_{0}^{\prime}(\sigma)\sigma=\int_{\{\sigma(x)>0\}}E_{0}^{\prime}(\sigma)\sigma=\lambda_{1}\int_{\{\sigma(x)>0\}}\sigma=\lambda_{1}$
@@
$\int E_{0}^{\prime}(\sigma)\sigma=\int\left(A^{\prime}(\sigma)-V_{\sigma}\right)\sigma=\gamma U(\sigma)-G(\sigma,\sigma)$
@@ 2.3 Uniqueness Results for Non-rotating Bodies
In this subsection, we discuss uniqueness results. To show the uniqueness result in the following, we follow Lieb and Yau’s arguments [13, Lemma 10, Lemma 11 and remark following] [16], but adapt it from the framework of quantum mechanics to that of classical mechanics. We show if, in addition, $P(\rho)$ satisfies (F4), $A^{\prime}\left(\rho^{3}\right)$ has second derivative and is convex, uniqueness of minimizer up to translation can also be shown in the following arguments. In particular, if polytropic law holds true with $\gamma>\frac{4}{3}$, then $A^{\prime}\left(\rho^{3}\right)=\frac{K\gamma}{\gamma-1}\rho^{3(\gamma-1)}$ satisfies those conditions.
Due to Theorem 2.6 (iii), we know the minimizer $\sigma_{m}$ is spherically symmetric after translation. In general, given a spherically symmetric function $\rho$, with an abuse of notation, we denote $\rho(|x|)=\rho(s),m(r)=m_{\rho}(r)=\int_{|x|<r}\rho=4\pi\int_{0}^{r}\rho(s)s^{2}ds$ is mass of $\rho$ in $\{|x|<r\}$, then easy to show (we assume the following computational results are finite. This condition holds, for instance, when $\rho$ is continuous with compact support):
@@
By directly computation we have
@@
Note that $V_{\rho}(0)=V_{\rho}(|0|)=4\pi\int_{0}^{\infty}t\rho(t)dt$, which follows directly from the definition of $V_{\rho}$ (or by applying L’Hôpital’s rule to show $V_{\rho}(0)=\lim\limits_{x\rightarrow 0}V_{\rho}(x)=4\pi\int_{0}^{\infty}t\rho(t)dt$). Then after using L’Hôpital’s rule [20, 5.13 Theorem], one can also show that
@@
Another way to show $\nabla V_{\rho}(0)=0$ is the fact that since $\lim\limits_{x\rightarrow 0}\nabla V_{\rho}(x)=0$ exists, by mean value theorem we can prove $\nabla V_{\rho}(0)=\lim\limits_{x\rightarrow 0}\nabla V_{\rho}(x)=0$. The third way to show it is using the fact $V_{\rho}$ is also spherically symmetric (same for $\frac{d}{dr}V_{\rho}(0)$) [^5] Strictly speaking, we need to verify that $V$ is differentiable at 0 then use symmetric property. This differentiability can be deduced from the regularity of $\rho$. For instance, if $\rho$ is continuous with compact support, then arguments similar to those in Theorem 2.6 (vi) ensure the differentiability of $V$ at 0..
@@
Theorem 2.31 (Shell Theorem [22] [19, Page 192 Theorem XXX and Page 193 Theorem XXXI] [18, Page 218 Theorem XXX and Theorem XXXI]).
@@
Without loss of generality, we can set up the coordinate system such that the center of $\rho$ is $(0,0,0)$, and the center of $\sigma$ is $\mathbf{D}=(0,0,R)$. Assume the radius of $\rho$ is $r_{1}$ and the radius of $\sigma$ is $r_{2}$, notice $R\geq r_{1}+r_{2}$, then thanks to Theorem 2.31 the mutual gravitational potential energy is given by:
@@
Let’s recall that for minimizer $\sigma$ (not necessary with mass 1 here), by Theorem 2.6 (vii) we have Euler-Lagrange equation (EL):
@@
On the other hand, thanks to Lieb and Yau [13, Lemma 8], we can show $0<R(\beta)<\infty$ holds true for any radial solutions to (38) under certain assumptions of initial conditions mentioned above and of pressure [13]. Therefore, the positive solutions to (38) can correspond one-to-one with the solutions to (EL). We will return to discuss more about these relations later.
@@
Lemma 2.36 (Relation between Central Density and Mass [13, Lemma 10]).
Suppose $\sigma_{m}$ and $\sigma_{\widetilde{m}}$ are minimizers for $E_{0}(\rho)$ with $\int\sigma_{m}=m,\int\sigma_{\widetilde{m}}=\widetilde{m}$ respectively and their centers of mass are $(0,0,0)^{T}$. Let $R_{m}$ and $R_{\widetilde{m}}$ be the radii of their supports and set $R=\max\left\{R_{m},R_{\widetilde{m}}\right\}$. Suppose that $\sigma_{m}(0)>\sigma_{\widetilde{m}}(0)$, then $\forall 0<r<R$ we have $m(r)>\widetilde{m}(r)$.
@@
We consider dividing the density functions into inner parts and outer parts. That is, let $\sigma_{m}^{i}=\sigma_{m}\mathbf{1}_{\{|x|\leq\bar{r}\}},\sigma_{m}^{o}=\sigma_{m}\mathbf{1}_{\{|x|>\bar{r}\}},\sigma_{\widetilde{m}}^{i}=\sigma_{\widetilde{m}}\mathbf{1}_{\{|x|\leq\bar{r}\}},\sigma_{\widetilde{m}}^{o}=\sigma_{\widetilde{m}}\mathbf{1}_{\{|x|>\bar{r}\}}$. Set $E_{0}^{i}\left(\rho^{i}\right)=E_{0}\left(\rho^{i}\right)$, and
$E_{0}^{o}\left(\rho^{o}\right)=U\left(\rho^{o}\right)-\frac{G\left(\rho^{o},\rho^{o}\right)}{2}-Q\int\frac{\rho^{o}(x)}{|x|}d^{3}x$
@@
$U^{i}=\left\{\rho^{i}\in L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)\middle|\rho^{i}\geq 0,\rho^{i}\text{ vanishes outside }\overset{-}{\left({B_{\widetilde{r}}(0)}\right)},~{\int\rho^{i}}=Q\right\}$
@@
Notice that in Lemma 2.36, only when $\forall 0<r<R$ we have $m(r)>\widetilde{m}(r)$. Hence it does not say $m>\widetilde{m}$, although we shall later see that this is true, for example due to Corollary 2.44 or Remark 3.3, but we do not yet know it. If we knew in advance that $m>\widetilde{m}$, the proof of the uniqueness of minimizer would be trivial.
It’s conceivable that the structure of pressure $P(\rho)$, and thereby $A(\rho)$ would play a role in finding uniqueness results. Previously, we only assume $P(\rho)$ satisfies assumptions (F1)(F2)(F3), but later we may need to assume (F4) also holds since we may take into account some derivatives.
However, so far, we only know the energy consists of inertial energy and gravitational interaction energy:
@@
To find a deeper relation between energy and $A(\rho)$, it will be nice if we can find another expression of $E_{0}(\rho)$ which $G(\rho,\rho)$ does not appear. Thanks to Lieb and Yau’, arguments [13], we can truly find such expression when $\rho$ is a minimizer, since in this case we have Euler-Lagrange equations which connects $A^{\prime}(\rho)$ with gravitational potential $V_{\rho}$ and thereby $G(\rho,\rho)$.
Lemma 2.38 (Relation between $E_{0}$ and $A$ [13, Lemma 6]).
Let $P(s)$ satisfies (F1) (F2) (F3). Let $\sigma$ is continuous, spherically symmetric with finite mass $m=\int\sigma,\sigma$ has compact support, and satisfies (EL) on all of $\mathbb{R}^{3}$ for $J=0$ and a single $\lambda<0$. Then
$E_{0}(\sigma)=\int\left(4A(\sigma)-3\sigma A^{\prime}(\sigma)\right)$ (40)

### v2

@@ 2.2 Existence Results for Non-rotating Bodies
$=f(r,\pi)$
Since $r\geq r_{0}$, when $\pi$ is large enough we have $f(r,\pi)\leq-l$ as we mention above. Therefore, the claim holds true. Thus $\forall\pi>\pi_{0}$, $\forall x$ with $r(x)>\pi$, we have $V_{\sigma_{R}}(x)\leq f(\frac{2r_{0}+1+\pi}{4},\pi)\leq-l$, then $\sigma_{R}(x)=0$ due to (23) and Lemma 2.24.
Now we consider the z-direction and show $V_{\sigma_{R}}(x)\leq-l$ holds when $z(x)$ is large enough, which is relatively easier. We can first show $\exists\delta>0$, $r\geq 2$ and $\widetilde{R_{0}}>0$, such that $\forall R\geq\widetilde{R_{0}}$, if $\int_{|z(x)-d|<r}\sigma_{R}(x)\,dx<\delta$, then $\sigma_{R}(x)=0$ for $|z(x)-d|\leq 1$. This comes from that when $|z(x)-d|\leq 1$ we have the estimate:
$V_{\sigma_{R}}(x)=\int_{|z(y)-d|<r}\frac{\sigma_{R}(x)}{|y-x|}\,dy+\int_{|z(y)-d|>r}\frac{\sigma_{R}(x)}{|y-x|}\,dy\leq c\left(\delta^{a}+\delta^{b}\right)+\frac{m}{r-1}$
And then letting $\delta$ small enough and $r$ large enough can give $V_{\sigma_{R}}(x)\leq-l$, which means $\sigma_{R}(x)=0$. Then we set $Z_{n}=\{x||z(x)-2n\mid\leq 1\},n=0,\pm 1,\pm 2,\ldots$. and $Z_{n}^{\prime}=\{x||z(x)-2n\mid<r\}$, without loss of generality, we can take $r$ to be an integer. Let $N$ be the number of $Z_{n}^{\prime}$ such that $\int_{Z_{n}^{\prime}}\sigma_{R}\,dx\geq\delta$. Since the mass of $\sigma_{R}$ is $m<\infty$ and the $Z_{n}^{\prime}$s overlap finite times $\widetilde{r}$$\infty$, we have $N\delta\leq\widetilde{r}m$, i.e., $N\leq\frac{\widetilde{r}m}{\delta}$. Therefore, $\sigma_{R}(x)=0$ for large $R$ except on a set of horizontal slabs $Z_{n}$ of total height at most $\frac{2\widetilde{r}m}{\delta}$.
If these $Z_{n}{s}$ were known to be contiguous, then we would have $\sigma_{R}(x)=0$ for $|z(x)|\geq\frac{2rm}{\delta}$. But if the $Z_{n}{s}$ were not contiguous, we could follow the idea of the strong rearrangement inequality in Lieb [43, Lemma 3]. We could rearrange the order of $Z_{n}$, slide the half-space above an empty $Z_{n}$ down one unit and obtain a new $\widetilde{\sigma_{R}}\in W_{R}$ from $\sigma_{R}$. During the process we notice
$U\left(\widetilde{\sigma_{R}}\right)=\int_{\mathbb{R}^{3}}A\left(\widetilde{\sigma_{R}}\right)\,dx=\int_{\mathbb{R}^{3}}A\left(\sigma_{R}\right)\,dx=U\left(\sigma_{R}\right)$
$G\left(\widetilde{\sigma_{R}},\widetilde{\sigma_{R}}\right)=\int_{\mathbb{R}^{3}}\widetilde{\sigma_{R}}V_{\widetilde{\sigma_{R}}}\,dx>\int_{\mathbb{R}^{3}}\sigma_{R}V_{\sigma_{R}}\,dx=G\left(\sigma_{R},\sigma_{R}\right)$
then we obtain $E_{0}\left(\widetilde{\sigma_{R}}\right)<E_{0}\left(\sigma_{R}\right)$, which leads to a contradiction.
@@
Thanks to Lemma 2.21 and Lemma 2.27, we know $\forall R\geq R^{\prime},\sigma_{R}\in W_{R^{\prime}}$, which implies $E_{0}\left(\sigma_{R}\right)=E_{0}\left(\sigma_{R^{\prime}}\right)$ since they should both be minimizers on $W_{R^{\prime}}$. Thus, we just need to show $\sigma_{R^{\prime}}$ is a minimizer on $W_{\infty}=m{R}\left(\mathbb{R}^{3}\right)$. To show that, given $\rho\in W_{\infty}$, we construct a $\rho_{R}\in W_{R}$ for large $R$, with $E_{0}\left(\rho_{R}\right)\leq E_{0}(\rho)+\epsilon$, where $\epsilon$ can be arbitrarily small. (For the construction of $\rho_{R}$, although $\rho$ is not necessarily to be spherically symmetric, but one can apply a strong rearrangement inequality in Lieb [43, Lemma 3] and the generalization by Sobolev of the Hardy-Littlewood theorem on rearrangements of functions [54] to replace $\rho$ by another spherically symmetric and radially decreasing (after translation) $\rho$, see for example [4, Appendix], [46, Section 6]. Then one can refer to [4, Section 6] or [16, Section 4] to construct $\rho_{R}$.) Then we have $E_{0}\left(\sigma_{R^{\prime}}\right)=E_{0}\left(\sigma_{R}\right)\leq E_{0}\left(\rho_{R}\right)\leq E_{0}(\rho)+\epsilon$. Let $\epsilon\rightarrow 0$ we obtain $E_{0}\left(\sigma_{R^{\prime}}\right)\leq E_{0}(\rho)$, which shows the result. ∎
It is natural to conjecture whether there are some relations between Lagrange-multipliers $\lambda_{m}$ and the derivatives of $e_{0}(m)$, since they can be obtained by somehow “differentiating” $E_{0}\left(\sigma_{m}\right)$. Thanks to Lieb and Yau’s paper [42, Theorem 3 and Theorem 4], we can know the left and right derivatives of $e_{0}(m)$ bound $\lambda:e_{0}^{\prime}\left(m^{+}\right)\leq\lambda\leq e_{0}^{\prime}\left(m^{-}\right)$, as cited in [46, Theorem 3.5] and [17, Section 2]. Moreover, we find a more precise quantitative property in the case that complies with polytropic equations of state in the following.
@@
By scaling relations in Section 3 (Theorem 3.2) we can obtain $e_{0}(m)=m^{\frac{5\gamma-6}{3\gamma-4}}e_{0}$, which implies $e_{0}^{\prime}$ exists. Then $e_{0}^{\prime}(m)=\lambda_{m}$ can come from Theorem 2.6 (viii). As in the proof of Theorem 2.6 (ii), we use the scaling relations mentioned in Section 3 (Proposition 3.4). We obtain $E_{0}^{\prime}\left(\sigma_{m}\right)(x)=\frac{1}{A^{\gamma-1}}E_{0}^{\prime}(\sigma)\left(\frac{1}{B}x\right)$, thus $\lambda_{m}=\frac{1}{A^{\gamma-1}}\lambda_{1}$. Since $e_{0}(m)=m^{\frac{5\gamma-6}{3\gamma-4}}e_{0}$, then $e_{0}^{\prime}(m)=\frac{5\gamma-6}{3\gamma-4}m^{\frac{2\gamma-2}{3\gamma-4}}e_{0}$, and $e_{0}^{\prime}(m)=\frac{1}{A^{\gamma-1}}e_{0}^{\prime}(1)$. Therefore, we just need to consider the case $m=1$. Notice we have Euler-Lagrange equation $E_{0}^{\prime}(\sigma)(x)=\lambda_{1}$ a.e. on $\{\sigma(x)>0\}$. Rather than substituting a specific $x$ to calculate $E_{0}^{\prime}(\sigma)$, we view $E_{0}^{\prime}(\sigma)$ as a functional in the sense that $E_{0}^{\prime}(\sigma)(\rho)=\int_{\mathbb{R}^{3}}E_{0}^{\prime}(\sigma)\rho\,dx$ as in Lemma 2.9 or Remark 2.10. We take $\rho=\sigma$ as a test function, notice the mass of $\sigma$ is 1. Then we have
$\int_{\mathbb{R}^{3}}E_{0}^{\prime}(\sigma)\sigma\,dx=\int_{\{\sigma(x)>0\}}E_{0}^{\prime}(\sigma)\sigma\,dx=\lambda_{1}\int_{\{\sigma(x)>0\}}\sigma\,dx=\lambda_{1}$
@@
$\int_{\mathbb{R}^{3}}E_{0}^{\prime}(\sigma)\sigma\,dx=\int_{\mathbb{R}^{3}}\left(A^{\prime}(\sigma)-V_{\sigma}\right)\sigma\,dx=\gamma U(\sigma)-G(\sigma,\sigma)$
@@ 2.3 Uniqueness Results for Non-rotating Bodies
In this subsection, we discuss uniqueness results. To show the uniqueness result in the following, we follow Lieb and Yau’s arguments [42, Lemma 10, Lemma 11 and remark following] [46], but adapt it from the framework of quantum mechanics to that of classical mechanics. We show if, in addition, $P(\rho)$ satisfies (F4), $A^{\prime}\left(\rho^{3}\right)$ has second derivative and is convex, uniqueness of minimizer up to translation can also be shown in the following arguments. In particular, if polytropic law holds true with $\gamma>\frac{4}{3}$, then $A^{\prime}\left(\rho^{3}\right)=\frac{K\gamma}{\gamma-1}\rho^{3(\gamma-1)}$ satisfies those conditions.
Due to Theorem 2.6 (iii), we know the minimizer $\sigma_{m}$ is spherically symmetric after translation. In general, given a spherically symmetric function $\rho$, with an abuse of notation, we denote $\rho(|x|)=\rho(s)$, and
$m(r)=m_{\rho}(r)=\int_{|x|<r}\rho\,dx=4\pi\int_{0}^{r}\rho(s)s^{2}ds$
is the mass of $\rho$ in $\{|x|<r\}$, then easy to show (we assume the following computational results are finite. This condition holds, for instance, when $\rho$ is continuous with compact support):
@@
By direct computation we have
@@
Note that $V_{\rho}(0)=V_{\rho}(|0|)=4\pi\int_{0}^{\infty}t\rho(t)dt$, which follows directly from the definition of $V_{\rho}$ (or by applying L’Hôpital’s rule to show $V_{\rho}(0)=\lim\limits_{x\rightarrow 0}V_{\rho}(x)=4\pi\int_{0}^{\infty}t\rho(t)dt$). Then after using L’Hôpital’s rule [52, Theorem 5.13], one can also show that
@@
Another way to show $\nabla V_{\rho}(0)=0$ is the fact that since $\lim\limits_{x\rightarrow 0}\nabla V_{\rho}(x)=0$ exists, by mean value theorem we can prove $\nabla V_{\rho}(0)=\lim\limits_{x\rightarrow 0}\nabla V_{\rho}(x)=0$. The third way to show it is using the fact $V_{\rho}$ is also spherically symmetric (same for $\frac{d}{dr}V_{\rho}(0)$) [^5] Strictly speaking, we need to verify that $V$ is differentiable at 0 then use the symmetric property. This differentiability can be deduced from the regularity of $\rho$. For instance, if $\rho$ is continuous with compact support, then arguments similar to those in Theorem 2.6 (vi) ensure the differentiability of $V$ at 0..
@@
Theorem 2.30 (Shell Theorem [57] [49, Page 192 Theorem XXX and Page 193 Theorem XXXI] [48, Page 218 Theorem XXX and Theorem XXXI]).
@@
Without loss of generality, we can set up a coordinate system such that the center of $\rho$ is $(0,0,0)$, and the center of $\sigma$ is $\mathbf{D}=(0,0,R)$. Assume the radius of $\rho$ is $r_{1}$ and the radius of $\sigma$ is $r_{2}$, notice $R\geq r_{1}+r_{2}$, then thanks to Theorem 2.30 the mutual gravitational potential energy is given by:
@@
Let’s recall that for minimizer $\sigma$, by Theorem 2.6 (vii) we have the Euler-Lagrange equation (EL):
$A^{\prime}(\sigma(x))=\left[V_{\sigma}(x)+\lambda\right]_{+}$ (EL)
@@
On the other hand, thanks to Lieb and Yau [42, Lemma 8], we can show $0<R(\beta)<\infty$ holds true for any radial solutions to (37) under certain assumptions of initial conditions mentioned above and of pressure [42]. Furthermore, note that our preceding arguments are not restricted to a specific mass (e.g., $m=1$), but naturally generalizes to any mass $m>0$. Consequently, the positive solutions to (37) actually correspond one-to-one with the solutions to (EL). We will return to discuss more about these relations later.
@@
Lemma 2.35 (Relation between Central Density and Mass [42, Lemma 10]).
Suppose $\sigma_{m}$ and $\sigma_{\widetilde{m}}$ are minimizers for $E_{0}(\rho)$ with $\int_{\mathbb{R}^{3}}\sigma_{m}\,dx=m$, $\int_{\mathbb{R}^{3}}\sigma_{\widetilde{m}}\,dx=\widetilde{m}$ respectively and their centers of mass are $(0,0,0)^{T}$. Let $R_{m}$ and $R_{\widetilde{m}}$ be the radii of their supports and set $R=\max\left\{R_{m},R_{\widetilde{m}}\right\}$. Suppose that $\sigma_{m}(0)>\sigma_{\widetilde{m}}(0)$, then $\forall 0<r<R$ we have $m(r)>\widetilde{m}(r)$.
@@
We consider dividing the density functions into inner parts and outer parts. That is, let
$\sigma_{m}^{i}=\sigma_{m}\mathbf{1}_{\{|x|\leq\bar{r}\}},\quad\sigma_{m}^{o}=\sigma_{m}\mathbf{1}_{\{|x|>\bar{r}\}},$
$\sigma_{\widetilde{m}}^{i}=\sigma_{\widetilde{m}}\mathbf{1}_{\{|x|\leq\bar{r}\}},\quad\sigma_{\widetilde{m}}^{o}=\sigma_{\widetilde{m}}\mathbf{1}_{\{|x|>\bar{r}\}}.$
Set $E_{0}^{i}\left(\rho^{i}\right)=E_{0}\left(\rho^{i}\right)$, and
$E_{0}^{o}\left(\rho^{o}\right)=U\left(\rho^{o}\right)-\frac{G\left(\rho^{o},\rho^{o}\right)}{2}-Q\int_{\mathbb{R}^{3}}\frac{\rho^{o}(x)}{|x|}\,dx$
By Remark 2.30, $\int_{\mathbb{R}^{3}}\sigma_{m}^{i}\,dx=\int_{\{|x|\leq\widetilde{r}\}}\sigma_{m}^{i}\,dx=Q$, and notice
$G\left(\sigma_{m}^{i},\sigma_{m}^{o}\right)$ $=\iint_{\mathbb{R}^{3}\times\mathbb{R}^{3}}\frac{\sigma_{m}^{i}(y)\sigma_{m}^{o}(x)}{|x-y|}\,dxdy$
$=\int_{\{|y|\leq\widetilde{r}\}}\sigma_{m}^{i}(y)V_{\sigma_{m}^{o}}(y)\,dy$
@@
$=QV_{\sigma_{m}^{o}}(0)$
$=Q\int_{\mathbb{R}^{3}}\frac{\sigma_{m}^{o}(x)}{|x|}\,dx.$
we obtain $e_{0}(m)=E_{0}\left(\sigma_{m}\right)=E_{0}^{i}\left(\sigma_{m}^{i}\right)+E_{0}^{o}\left(\sigma_{m}^{o}\right)$. Similarly we have $e_{0}(\widetilde{m})=E_{0}\left(\sigma_{\widetilde{m}}\right)=E_{0}^{i}\left(\sigma_{m}^{i}\right)+E_{0}^{o}\left(\sigma_{\widetilde{m}}^{o}\right)$.
@@
$U^{i}=\left\{\rho^{i}\in L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)\middle|\rho^{i}\geq 0,\rho^{i}\text{ vanishes outside }\overline{B_{\widetilde{r}}(0)},~{\int_{\mathbb{R}^{3}}\rho^{i}\,dx}=Q\right\}$
@@
Notice that in Lemma 2.35, only when $\forall 0<r<R$ we have $m(r)>\widetilde{m}(r)$. Hence it does not say $m>\widetilde{m}$, although we shall later see that this is true (Corollary 2.43 or Remark 3.3) after obtaining the uniqueness result. If we knew in advance that $m>\widetilde{m}$, the proof of the uniqueness of minimizer would be trivial.
So far we only know the energy consists of inertial energy and gravitational interaction energy:
@@
To find a deeper relation between energy and $A(\rho)$, it will be nice if we can find another expression of $E_{0}(\rho)$ in which $G(\rho,\rho)$ does not appear. Thanks to Lieb and Yau’s arguments [42], we can truly find such expression when $\rho$ is a minimizer, since in this case we have Euler-Lagrange equation which connects $A^{\prime}(\rho)$ with gravitational potential $V_{\rho}$ and thereby $G(\rho,\rho)$.
Lemma 2.37 (Relation between $E_{0}$ and $A$ [42, Lemma 6]).
Let $P(s)$ satisfies (F1) (F2) (F3). Suppose $\sigma$ is continuous, spherically symmetric with finite mass $m=\int_{\mathbb{R}^{3}}\sigma\,dx$, $\sigma$ has compact support, and satisfies (EL) on all of $\mathbb{R}^{3}$ for $J=0$ and a single $\lambda<0$. Then
$E_{0}(\sigma)=\int_{\mathbb{R}^{3}}\left(4A(\sigma)-3\sigma A^{\prime}(\sigma)\right)\,dx$ (39)